# A Polarization Hall Effect in Hydrated DNA

### v1

@@ Abstract
Understanding how biological soft matter responds to electromagnetic fields under ambient conditions remains a central challenge, as thermal fluctuations are generally expected to suppress long-range organization. Here, we report that hydrated DNA exhibits a reproducible magnetic-field-induced transition characterized by a sharp transverse-voltage threshold (– mV), followed by a regime of regular, phase-stable oscillations in the transverse polarization signal. These features emerge only beyond the threshold and display a pronounced temperature dependence, consistent with the formation of a collective mode within the hydrogen-bond network of the DNA–water interface. Motivated by recent studies of Hall-like responses carried by neutral excitations, including phonons, magnons, and excitons, we interpret the observed transverse signal in terms of coherent polarization dynamics of proton–proton-hole dipoles confined to a quasi-two-dimensional hydrated layer. Within this framework, the transverse response is attributed to a field-organized polarization mode; the measured transverse voltage arises from collective dipolar dynamics rather than steady carrier transport. These results identify hydrated DNA as a soft-matter system in which magnetic field and temperature jointly modulate collective polarization dynamics, providing a biologically relevant platform for exploring coherence and transverse responses in hydrogen-bonded media.
@@ I Introduction
Hydrated DNA forms a structured yet dynamically fluctuating hydrogen-bond network whose collective electrodynamic properties remain only partially understood. At the molecular scale, proton transfer along hydrogen bonds, hydration-layer ordering, and long-range dipolar interactions have been studied extensively, with theoretical models predicting cooperative proton dynamics and correlated vibrational modes in soft biological assemblies [Frohlich1968, Frohlich1988, Davydov1979, HammesSchiffer2006, DonadioGalli2025]. These microscopic ideas are complemented by macroscopic frameworks developed to describe coherent or synergistic organization in biological systems [Reimers2009, Vasconcellos2012], as well as by broader discussions of biological coherence and quantum-like behavior [AlKhaliliMcFadden2015].
@@
A central open question is whether such collective behavior can persist at mesoscopic scales under ambient conditions, given the well-known constraints imposed by thermal fluctuations and decoherence in biological environments [Tegmark2000]. Most experimental and theoretical studies of DNA electrostatics have focused on ionic conduction in solution or charge migration along -stacked base pairs, leaving unresolved whether coherent polarization modes might instead arise within the DNA–water interface under external-field excitation. Prior analyses of cooperative protonic and dipolar dynamics in biological matter, together with atomistic simulations of confined hydration layers [DonadioGalli2025], suggest that a structured hydration shell could, in principle, support correlated dipolar fluctuations. Direct macroscopic evidence for such field-induced collective behavior, however, has remained limited.
@@
Over the past several years, magnetic-field experiments on hydrated DNA have revealed a reproducible pattern of sudden voltage thresholds, staircase-like plateaus, and oscillatory features at or near room temperature [PietruszkaMarzec2024, Pietruszka2025a, Pietruszka2025b]. The observed -periodic oscillations bear a formal resemblance to Shubnikov–de Haas behavior known from electronic quantum transport [Shoenberg1984, Tsui1982, VonKlitzing1980]. At the same time, the biological context and the absence of sustained charge flow point to a fundamentally different physical mechanism. Control measurements on pure water, desiccated DNA, and ionic or metallic substrates do not reproduce these effects, indicating that the structured DNA hydration layer—rather than conventional electrical conduction—governs the response.
@@
The physical interpretation of these features has therefore required reassessment. Earlier work described the transverse signals in terms of a “protonic Hall effect,” implicitly invoking mobile proton carriers. Subsequent experiments combining magnetic-field sweeps with high-resolution temperature control demonstrated that the transverse response persists even in regimes where no continuous conduction pathway exists. These observations motivated a reinterpretation in which the measured voltages arise from collective polarization dynamics of proton–proton-hole dipoles embedded within the hydrogen-bond network of the hydrated DNA interface.
@@
Importantly, neutral dipolar excitations can exhibit Hall-type transverse responses. In modern condensed-matter physics, it is now well established that neutral quasiparticles—including phonons, magnons, and excitons—can generate Hall-like effects through mechanisms involving Berry curvature, chiral geometry, or broken inversion symmetry [Strohm2005, Onose2010, Mak2018, Saito2019, Li2023, ZhangGaoChen2024, Jin2025, Owerre2016]. These developments provide a natural conceptual bridge: hydrated DNA hosts strongly polarizable protonic dipoles within an intrinsically chiral molecular environment, making a magnetically induced transverse polarization response plausible even in the absence of itinerant charge carriers.
@@
From this perspective, hydrated DNA can be viewed as a soft, quasi-two-dimensional dipolar medium whose polarization modes are organized jointly by magnetic field and temperature. The possibility of macroscopic coherence in such dipolar ensembles is consistent with early predictions for driven biological systems [Frohlich1968, Reimers2009] and with general principles of phase coherence and mode condensation studied in physical systems, including Bose–Einstein condensates [Anderson1995, Ketterle1999]. Although the DNA–water interface operates in a regime very different from ultracold atomic gases, the underlying concept—the emergence of collective phase coherence from coupled degrees of freedom—remains relevant.
@@
Here we employ a combined magnetic-field and temperature-control strategy to systematically map coherent regimes in hydrated DNA. By treating magnetic field and temperature as independent control parameters and by monitoring simultaneous longitudinal and transverse responses, we resolve transitions between incoherent fluctuations, SdH-like oscillatory behavior, plateau-like polarization states, and a low-temperature Fröhlich-type coherent mode. Together, these results demonstrate that hydrated DNA supports field-stabilized and temperature-gated polarization coherence, extending the physics of neutral-excitation Hall responses and coherent dipolar ordering to a biologically relevant, hydrogen-bonded soft-matter system.
@@ II Results
Under moderate magnetic fields and ambient conditions, we observe three canonical features in this polarization channel: (i) -periodic polarization oscillations in both and , which act as SdH-like precursors of dipolar quantization (used here as a transport-style formalism rather than evidence of electronic Landau quantization); (ii) a sharp, field-driven threshold marking reorganization of dipolar domains; and (iii) discrete, Landau-like polarization plateaus in .
Both the oscillatory periodicity and the plateau spacing yield convergent estimates of an effective coherence density, , indicating that a single field-organized dipolar ensemble likely governs the response. A transition near T marks a reorganization from repeated Landau-like dipolar quantization to a Fröhlich-like phase characterized by long-range, phase-locked dipolar oscillations and simplified response lobes. Control measurements in pure water lack these features, confirming that structured DNA hydration layers are essential.
Throughout this work we retain the quantitative transport-style formalism (e.g., periodicity, filling-factor-style plateaus, and tensor inversion), but interpret the observables within a polarization-response framework. In this view, extracted scales (e.g., densities inferred from periodicities or plateau spacings) parameterize the collective dynamics of a field-organized dipolar ensemble rather than itinerant charge transport. The hydrated DNA interface thus operates as a quasi-two-dimensional dipolar medium in which the magnetic field discretizes polarization modes and stabilizes macroscopic dipolar coherence at room temperature.
@@
To probe field-induced coherence, we measured the transverse polarization voltage in a quasi-two-dimensional hydrated DNA layer under a perpendicular magnetic field at ambient conditions. Figure 1 provides the microscopic basis adopted throughout: transient proton relays along hydrogen bonds generate locally neutral proton–proton-hole dipoles that align and oscillate cooperatively within the hydration layer. These dipoles experience a Lorentz-like magnetic torque, and their collective reorientation generates a measurable polarization current , detected macroscopically as the transverse voltage . Under moderate magnetic fields (–) and room temperature, this polarization channel exhibits a reproducible progression of features: (i) -periodic oscillations in both and , serving as Shubnikov–de Haas-like precursors of dipolar quantization; (ii) a sharp, field-driven threshold; and (iii) the emergence of discrete, Landau-like polarization plateaus in . The oscillation periodicity and the spacing of these plateaus both yield a convergent coherence density of approximately , indicating that a single, field-organized dipolar ensemble governs the response. Control measurements in pure water remain featureless under otherwise identical conditions, demonstrating that structured DNA hydration layers are essential for the observed field-induced coherence.
@@
To resolve how temperature influences the organization of the dipolar network, we performed continuous cooling sweeps at a fixed magnetic field and monitored both longitudinal (and ) and transverse () polarization channels at 1 Hz sampling. Figures 2 and 3 summarize two complementary experiments performed on consecutive days using 500 ng/µL and 100 ng/µL DNA–water samples, respectively. In Figure 2, the longitudinal channels reveal two clear temperature-driven reorganizations at 20.6 °C and 12.0 °C, expressed as sharp changes in the fluctuation amplitude and correlation structure of and . In Figure 3, an independent sweep under identical magnetic-field conditions () shows that the transverse channel remains quiet above about 12 °C, but below this same temperature, the system reorganizes into a regime of large-amplitude, nearly periodic oscillations exceeding 150 mV. Together, these measurements establish 12 °C as a robust temperature boundary where the dipolar ensemble transitions into a strongly coherent, oscillatory state.
@@
The combined behavior of Figures 2 and 3 reveals a clear, two-stage temperature dependence of coherence in hydrated DNA. The first restructuring at 20.6 °C, visible only in the longitudinal channels (Fig. 2), marks the onset of enhanced dipolar sensitivity, where thermal agitation becomes sufficiently reduced for correlated hydrogen-bond reorientation to begin. However, the transverse channel remains largely featureless above 12 °C (Fig. 3), indicating that this initial transition does not yet produce a robust, phase-locked polarization mode.
The second transition at 12.0 °C is qualitatively different. Below this temperature, both experiments show a pronounced reorganization: in Figure 2, the longitudinal fluctuations collapse into a low-temperature branch, while in Figure 3 the transverse channel suddenly acquires large-amplitude, nearly periodic oscillations that persist for several degrees of cooling. The simultaneous change in both longitudinal and transverse responses demonstrates that the dipolar network undergoes a collective mode-selection process at 12 °C, entering a regime where cross-channel coherence becomes strong, and the transverse polarization mode dominates the macroscopic response.
@@
This two-stage progression – initial softening at 20.6 °C followed by full coherence below 12.0 °C – supports the interpretation that temperature tunes the coupling strength and phase ordering of the proton–proton-hole dipolar ensemble. Above 20.6 °C, dipoles fluctuate largely independently; between 20.6 and 12.0 °C, they begin reorganizing but remain only partially correlated; and below 12.0 °C, the system crosses a coherence boundary and forms a stable, collectively oscillating polarization mode. These temperature-dependent reorganizations parallel the field-induced transitions observed elsewhere in the manuscript, reinforcing the view that coherent polarization in hydrated DNA emerges when both magnetic alignment and thermal suppression of decoherence act together to stabilize collective dipolar motion.
@@
Having established the temperature boundaries for coherence (Figs. 2 and 3), we next examine how the transverse polarization responds to magnetic field at fixed, near-room temperature. At a DNA concentration of 500 ng L-1 (Fig. 4), exhibits a sharp threshold at , followed by a staircase of quantized polarization plateaus. The regular spacing of the plateau centers (linear fit ) persists despite thermal noise of approximately mV, indicating discrete, field-stabilized configurations of the dipolar network. Control measurements performed on pure water remain featureless under otherwise identical conditions, confirming that quantization requires the DNA-templated hydration architecture. The plateau spacing, together with the oscillation periodicity analyzed below, yields a consistent coherence density of approximately , interpreted as the areal density of coherently active dipolar domains. This value reappears independently in the Fourier analysis of the oscillatory regime and in the convergence analysis, supporting a single underlying coherence ensemble. A notable feature of the transverse voltage staircase is the presence of pronounced fluctuations localized at the boundaries between neighboring plateaus. These fluctuations are not uniformly distributed across the signal but appear predominantly in narrow intervals preceding each step. Despite their strength, the system reproducibly settles into a subsequent plateau, indicating that the staircase structure is robust against these fluctuations. Such behavior is characteristic of driven systems undergoing transitions between metastable collective states. Near a stability boundary, competing polarization configurations may become nearly degenerate, leading to enhanced fluctuations as the system explores multiple configurations. Once a new configuration becomes energetically favorable, the system rapidly relaxes into a stable state, giving rise to the next plateau. If the observed steps were dominated by stochastic noise or instrumental artefacts, one would expect the fluctuations to smear out the plateau structure or lead to irregular step spacing. Instead, the fluctuations are confined to the transition regions, while the plateaus themselves remain well defined. This behavior suggests that the fluctuations play an active role in mediating transitions between discrete states rather than obscuring an underlying continuous response. Similar fluctuation-assisted switching phenomena are widely observed in nonlinear and self-organized systems, where noise or thermal agitation enables transitions between neighboring attractors in a multistable landscape.
@@
Lowering the temperature strengthens phase locking among dipoles and reveals SdH-like, -periodic signatures preceding plateau formation. For a DNA concentration of 100 ng L-1 at 5 (Fig. 5), both and exhibit robust oscillations periodic in over the field range 0.13–0.30 T. Identical oscillation frequencies in the two channels ( T and T; T-1) indicate a common quantization origin, here attributed to discrete reorganizations of the coherent polarization state rather than charge orbits. FFT-inferred parameters ( and , as reported in Fig. 5) quantify the characteristic scale of coherence patterning within the hydration sheet. All 5 data were acquired in the same measurement geometry; for reference, the plateaued trace shown in Fig. 4 was recorded at 22.3 .
At 6 and 500 ng L-1 (Fig. 6), the background-subtracted transverse signal exhibits dense, nearly periodic modulation with a mean spacing of approximately 0.83 T-1 and a dominant FFT peak at T. Interpreted within the polarization-quantization formalism using an Onsager-type geometric estimator, this yields an effective coherence density of , consistent with the plateau-derived value obtained near room temperature. The transverse oscillations exceed the longitudinal response by roughly four orders of magnitude and are approximately out of phase with , as expected for an intrinsically transverse polarization mode with weak dissipation. For T, the oscillatory pattern simplifies into broader lobes, indicating a field-driven reorganization into a more globally coherent state.
@@
Converting and to and , followed by tensor inversion, yields the polarization-response tensor describing how the longitudinal and transverse channels co-vary under an applied magnetic field (Fig. 7). Both components trace smooth, continuous trajectories rather than exhibiting random scatter, indicating organized dynamics rather than noise-dominated behavior. Above the threshold field, the transverse component becomes dominant and the response evolves into broader, slowly varying lobes, signaling a crossover into a more globally phase-locked dipolar state. This crossover is observed reproducibly under both isothermal conditions and continuous magnetic-field sweeps, supporting its interpretation as an intrinsic feature of the DNA hydration architecture rather than a protocol-dependent effect.
@@
Control measurements performed on distilled water (Fig. 8a) show no oscillations, thresholds, or plateau structure, whereas DNA solutions at 1000 and 500 ng L-1 (Figs. 8b and 8c) exhibit the characteristic double-peak motif followed by discrete polarization plateaus. These features are reproducible in both magnetic-field sweep directions (arrows), indicating sweep-direction independence and the absence of hysteresis. Together, these comparisons establish that structured DNA hydration layers are required for the emergence of coherent polarization dynamics.
To assess the role of ionic contributions, we examined unbuffered DNA dissolved directly in distilled water (Fig. 9). Both the 100 and 480 ng L-1 samples display a pronounced threshold near T (voltage drop mV), followed by stabilization into Landau-like polarization levels with a spacing of approximately 1 meV, close to the meV value reported previously. This agreement indicates that the observed quantization is intrinsic to the hydrated DNA matrix and does not depend on buffer ions. A Cu-coated control sample exhibits a transverse response of opposite sign, consistent with electronic conduction, further distinguishing the polarization-driven response observed in DNA solutions.
@@
Beyond the quantized steps and -periodic oscillations, coherent coupling between the longitudinal and transverse polarization channels is directly revealed by the interference reconstruction shown in Fig. 10. By combining the simultaneously acquired signals and during a linear magnetic-field ramp, the interference map exhibits alternating constructive and destructive bands whose spacing and contrast remain stable across the entire 100–200 mT interval at 10 for 100 ng L-1 DNA. The region of maximal fringe intensity occurs near , consistent with the independently measured –phase lag and with Lissajous traces obtained under identical conditions (not shown). The persistence and sharpness of these interference fringes indicate that the two channels share a common, phase-coherent polarization mode. In this regime, the system behaves as a coupled dipolar oscillator with a well-defined phase relationship between its longitudinal and transverse components, reinforcing the interpretation that the transverse response originates from coherent polarization dynamics within the DNA–water hydrogen-bond network.
@@
The approximate magnetic-field–temperature (–) phase diagram shown in Fig. 13 summarizes the regimes of coherent polarization identified across all measurements performed to date. The diagram incorporates magnetic-field thresholds (0.25–0.35 T), two reproducible temperature transitions at 20.6 and 12.0 , polarization staircases, bands of SdH-like oscillations, and the emergence of low-temperature Fröhlich-type oscillations. Together, these observations delineate three principal domains: (i) a high-temperature incoherent regime without quantization; (ii) an intermediate regime in which discrete polarization plateaus coexist with -periodic oscillations; and (iii) a low-temperature, high-field regime characterized by a globally phase-locked polarization mode. Gaussian blurring was applied during construction to avoid artificially sharp boundaries, reflecting the intrinsic softness and heterogeneity of the hydrogen-bond network. The resulting diagram illustrates how magnetic alignment and thermal suppression of decoherence jointly organize the DNA–water matrix into quantized, and ultimately coherent, polarization states.

### v2

@@ Abstract
Understanding how soft matter systems, including biological ones, can develop collective electromagnetic phenomena under external fields at ambient conditions remains a central challenge, as thermal fluctuations are generally expected to suppress long-range organization. Here, we report that hydrated DNA exhibits a reproducible magnetic-field-induced transition characterized by a sharp transverse-voltage threshold, followed by a regime of regular, phase-stable oscillations in the transverse polarization signal. These features emerge only beyond the threshold and display a pronounced temperature dependence, consistent with the formation of a collective mode within the hydrogen-bond network of the DNA–water interface. Motivated by recent studies of Hall-like responses carried by neutral excitations, including phonons, magnons, and excitons, we interpret the observed transverse signal in terms of coherent polarization dynamics of proton–proton-hole dipoles confined to a quasi-two-dimensional hydrated layer. Within this framework, the transverse response is attributed to a field-organized polarization mode; the measured transverse voltage arises from collective dipolar dynamics rather than steady carrier transport. These results identify hydrated DNA as a soft-matter system in which magnetic field and temperature jointly modulate collective polarization dynamics, providing a biologically relevant platform for exploring coherence and transverse phenomena in hydrogen-bonded media.
@@ I Introduction
Hydrated DNA forms a structured yet dynamically fluctuating hydrogen-bond network whose collective electrodynamic properties remain only partially understood. At the molecular scale, proton transfer along hydrogen bonds, hydration-layer ordering, and long-range dipolar interactions have been studied extensively, with theoretical models predicting cooperative proton dynamics and correlated vibrational modes in soft biological assemblies [7, 8, 3, 10, 4]. These microscopic ideas are complemented by macroscopic frameworks developed to describe coherent or synergistic organization in biological systems [22, 28], as well as by broader discussions of biological coherence and quantum-like behavior [1].
@@
A central open question is whether such collective behavior can persist at mesoscopic scales under ambient conditions, given the well-known constraints imposed by thermal fluctuations and decoherence in biological environments [26]. Most experimental and theoretical studies of DNA electrostatics have focused on ionic conduction in solution or charge migration along -stacked base pairs, leaving unresolved whether coherent polarization modes might instead arise within the DNA–water interface under external-field excitation. Prior analyses of cooperative protonic and dipolar dynamics in biological matter, together with atomistic simulations of confined hydration layers [4], suggest that a structured hydration shell could, in principle, support correlated dipolar fluctuations. Direct macroscopic evidence for such field-induced collective behavior, however, has remained limited.
@@
Over the past several years, magnetic-field experiments on hydrated DNA have revealed a reproducible pattern of sudden voltage thresholds, staircase-like plateaus, and oscillatory features at or near room temperature [19, 21, 20]. The observed -periodic oscillations bear a formal resemblance to Shubnikov–de Haas behavior known from electronic quantum transport [24, 27, 29]. At the same time, the biological context and the absence of sustained charge flow point to a fundamentally different physical mechanism. Control measurements on pure water, desiccated DNA, and ionic or metallic substrates do not reproduce these effects, indicating that the structured DNA hydration layer—rather than conventional electrical conduction—governs the response.
@@
The physical interpretation of these features has therefore required reassessment. Earlier work described the transverse signals in terms of a “protonic Hall effect,” implicitly invoking mobile proton carriers. Subsequent experiments combining magnetic-field sweeps with high-resolution temperature control demonstrated that the transverse response persists even in regimes where no continuous conduction pathway exists. These observations motivated a reinterpretation in which the measured voltages arise from collective polarization dynamics of proton–proton-hole dipoles embedded within the hydrogen-bond network of the hydrated DNA interface. A closely related ohmic–polarization crossover, identified through independent longitudinal measurements, has recently been reported in hydrated DNA under comparable experimental conditions [18].
@@
Importantly, neutral dipolar excitations can exhibit Hall-type transverse responses. In modern condensed-matter physics, it is now well established that neutral quasiparticles—including phonons, magnons, and excitons—can generate Hall-like effects through mechanisms involving Berry curvature, chiral geometry, or broken inversion symmetry [25, 15, 14, 23, 13, 32, 11, 16]. These developments provide a natural conceptual bridge: hydrated DNA hosts strongly polarizable protonic dipoles within an intrinsically chiral molecular environment, making a magnetically induced transverse polarization response plausible even in the absence of itinerant charge carriers.
@@
From this perspective, hydrated DNA can be viewed as a soft, quasi-two-dimensional dipolar medium whose polarization modes are organized jointly by magnetic field and temperature. The possibility of macroscopic coherence in such dipolar ensembles is consistent with early predictions for driven biological systems [7, 22] and with general principles of phase coherence and mode condensation studied in physical systems, including Bose–Einstein condensates [2, 12]. Although the DNA–water interface operates in a regime very different from ultracold atomic gases, the underlying concept—the emergence of collective phase coherence from coupled degrees of freedom—remains relevant.
@@
Here we employ a combined magnetic-field and temperature-control strategy to systematically map coherent regimes in hydrated DNA. By treating magnetic field and temperature as independent control parameters and by monitoring simultaneous longitudinal and transverse responses, we resolve transitions between incoherent fluctuations, -periodic oscillatory behavior, plateau-like polarization states, and a low-temperature Fröhlich-type coherent mode. Together, these results demonstrate that hydrated DNA supports field-stabilized and temperature-gated polarization coherence, extending the physics of neutral-excitation Hall responses and coherent dipolar ordering to a biologically relevant, hydrogen-bonded soft-matter system.
@@ II Results
Under moderate magnetic fields and ambient conditions, we observe three canonical features in this polarization channel: (i) -periodic polarization oscillations in both and , which act as precursors of dipolar quantization (used here as a transport-style formalism rather than evidence of electronic magnetotransport quantization); (ii) a sharp, field-driven threshold marking reorganization of dipolar domains; and (iii) discrete polarization plateaus in .
Both the oscillatory periodicity and the plateau spacing yield convergent estimates of an effective coherence density, , indicating that a single field-organized dipolar ensemble likely governs the response. A transition near T marks a reorganization from repeated field-discretized dipolar quantization to a Fröhlich-like phase characterized by long-range, phase-locked dipolar oscillations and simplified response lobes. Control measurements in pure water lack these features, confirming that structured DNA hydration layers are essential.
Throughout this work, we retain the quantitative transport-style formalism (e.g., periodicity, filling-factor-style plateaus, and tensor inversion), but interpret the observables within a polarization-response framework. In this view, extracted scales (e.g., densities inferred from periodicities or plateau spacings) parameterize the collective dynamics of a field-organized dipolar ensemble rather than itinerant charge transport. The hydrated DNA interface thus operates as a quasi-two-dimensional dipolar medium in which the magnetic field discretizes polarization modes and stabilizes macroscopic dipolar coherence at room temperature.
@@
To probe field-induced coherence, we measured the transverse polarization voltage in a quasi-two-dimensional hydrated DNA layer under a perpendicular magnetic field at ambient conditions. Figure 1 illustrates the microscopic picture adopted throughout this work: transient proton relays along hydrogen bonds are interpreted as generating locally neutral proton–proton-hole dipoles, which can align and oscillate cooperatively within the hydration layer. These dipoles can experience a Lorentz-like magnetic torque, and their collective reorientation generates a measurable polarization current , detected macroscopically as the transverse voltage . Under moderate magnetic fields (–) and room temperature, this polarization channel exhibits a reproducible progression of features: (i) -periodic oscillations in both and , serving as transport-style analogues of dipolar quantization; (ii) a sharp, field-driven threshold; and (iii) the emergence of discrete polarization plateaus in . The oscillation periodicity and the spacing of these plateaus both yield a convergent coherence density of approximately , indicating that a single, field-organized dipolar ensemble governs the response. Control measurements in pure water remain featureless under otherwise identical conditions, demonstrating that structured DNA hydration layers are essential for the observed field-induced coherence.
@@
To resolve how temperature influences the organization of the dipolar network, we performed continuous cooling sweeps at a fixed magnetic field and monitored both longitudinal (and ) and transverse () polarization channels at 1 Hz sampling. Figure 2 and Figure 3 summarize two complementary experiments performed on consecutive days using 500 ng/L and 100 ng/L DNA–water samples, respectively. In Figure 2, the longitudinal channels reveal two clear temperature-driven reorganizations at 20.6 and 12.0 , expressed as sharp changes in the fluctuation amplitude and correlation structure of and . In Figure 3, an independent sweep under identical magnetic-field conditions () shows that the transverse channel remains quiet above about 12 , but below this same temperature, the system reorganizes into a regime of large-amplitude, nearly periodic oscillations exceeding 150 mV. Together, these measurements establish 12 as a robust temperature boundary where the dipolar ensemble transitions into a strongly coherent, oscillatory state.
@@
The combined behavior of Figure 2 and Figure 3 reveals a clear, two-stage temperature dependence of coherence in hydrated DNA. The first restructuring at 20.6 , visible only in the longitudinal channels (Figure 2), marks the onset of enhanced dipolar sensitivity, where thermal agitation becomes sufficiently reduced for correlated hydrogen-bond reorientation to begin. However, the transverse channel remains largely featureless above 12 (Figure 3), indicating that this initial transition does not yet produce a robust, phase-locked polarization mode.
The second transition at 12.0 is qualitatively different. Below this temperature, both experiments show a pronounced reorganization: in Figure 2, the longitudinal fluctuations collapse into a low-temperature branch, while in Figure 3 the transverse channel suddenly acquires large-amplitude, nearly periodic oscillations that persist for several degrees of cooling. The simultaneous change in both longitudinal and transverse responses demonstrates that the dipolar network undergoes a collective mode-selection process at 12 , entering a regime where cross-channel coherence becomes strong, and the transverse polarization mode dominates the macroscopic response.
@@
This two-stage progression – initial softening at 20.6 followed by full coherence below 12.0 – supports the interpretation that temperature tunes the coupling strength and phase ordering of the proton–proton-hole dipolar ensemble. Above 20.6 , dipoles fluctuate largely independently; between 20.6 and 12.0 , they begin reorganizing but remain only partially correlated; and below 12.0 , the system crosses a coherence boundary and forms a stable, collectively oscillating polarization mode. These temperature-dependent reorganizations parallel the field-induced transitions observed elsewhere in the manuscript, reinforcing the view that coherent polarization in hydrated DNA emerges when both magnetic alignment and thermal suppression of decoherence act together to stabilize collective dipolar motion.
@@
Having established the temperature boundaries for coherence (Figs. 2 and 3), we next examine how the transverse polarization responds to magnetic field at fixed, near-room temperature. At a DNA concentration of 500 ng L-1 (Figure 4), exhibits a sharp threshold at , followed by a staircase of quantized polarization plateaus. The regular spacing of the plateau centers (linear fit ) persists despite thermal noise of approximately mV, indicating discrete, field-stabilized configurations of the dipolar network. Control measurements performed on pure water remain featureless under otherwise identical conditions, confirming that quantization requires the DNA-templated hydration architecture. The plateau spacing, together with the oscillation periodicity analyzed below, yields a consistent coherence density of approximately , interpreted as the areal density of coherently active dipolar domains. This value reappears independently in the Fourier analysis of the oscillatory regime and in the convergence analysis, supporting a single underlying coherence ensemble. A notable feature of the transverse voltage staircase is the presence of pronounced fluctuations localized at the boundaries between neighboring plateaus. These fluctuations are not uniformly distributed across the signal but appear predominantly in narrow intervals preceding each step. Despite their strength, the system reproducibly settles into a subsequent plateau, indicating that the staircase structure is robust against these fluctuations. Such behavior is characteristic of driven systems undergoing transitions between distinct collective polarization configurations. Near a stability boundary, competing polarization configurations may become nearly degenerate, leading to enhanced fluctuations as the system explores multiple configurations. Once a new configuration becomes energetically favorable, the system rapidly relaxes into a stable state, giving rise to the next plateau. If the observed steps were dominated by stochastic noise or instrumental artefacts, one would expect the fluctuations to smear out the plateau structure or lead to irregular step spacing. Instead, the fluctuations are confined to the transition regions, while the plateaus themselves remain well defined. This behavior suggests that the fluctuations play an active role in mediating transitions between discrete states rather than obscuring an underlying continuous response. Similar fluctuation-assisted switching phenomena are widely observed in nonlinear and self-organized systems, where noise or thermal agitation enables transitions between neighboring stable configurations.
@@
Lowering the temperature strengthens phase locking among dipoles and reveals -periodic signatures preceding plateau formation. For a DNA concentration of 100 ng L-1 at 5 (Figure 5), both and exhibit robust oscillations periodic in over the field range 0.13–0.30 T. Identical oscillation frequencies in the two channels ( T and T; T-1) indicate a common quantization origin, here attributed to discrete reorganizations of the coherent polarization state rather than charge orbits. FFT-inferred parameters ( and , as reported in Figure 5) quantify the characteristic scale of coherence patterning within the hydration sheet. All 5 data were acquired in the same measurement geometry; for reference, the plateaued trace shown in Figure 4 was recorded at 22.3 .
At 6 and 500 ng L-1 (Figure 6), the background-subtracted transverse signal exhibits dense, nearly periodic modulation with a mean spacing of approximately 0.83 T-1 and a dominant FFT peak at T. Interpreted within the polarization-quantization formalism using an Onsager-type geometric estimator, this yields an effective coherence density of , consistent with the plateau-derived value obtained near room temperature. The transverse oscillations exceed the longitudinal response by roughly four orders of magnitude and are approximately out of phase with , as expected for an intrinsically transverse polarization mode with weak dissipation. For T, the oscillatory pattern simplifies into broader lobes, indicating a field-driven reorganization into a more globally coherent state.
Although the -periodic oscillations preceding the threshold and the transverse oscillations could in principle be suspected as artifacts related to magnetic-field variation, this interpretation is disfavored by experiments in which the mechanical drive is completely switched off and the magnetic field is held constant at its maximum value, yet similarly dense and high-amplitude oscillations persist when the system is driven solely by slow temperature variation using passive cooling (ice).
@@
Converting and to and , followed by tensor inversion, yields the polarization-response tensor describing how the longitudinal and transverse channels co-vary under an applied magnetic field (Figure 7). Both components trace smooth, continuous trajectories rather than exhibiting random scatter, indicating organized dynamics rather than noise-dominated behavior. Above the threshold field, the transverse component becomes dominant and the response evolves into broader, slowly varying lobes, signaling a crossover into a more globally phase-locked dipolar state. This crossover is observed reproducibly under both isothermal conditions and continuous magnetic-field sweeps, supporting its interpretation as an intrinsic feature of the DNA hydration architecture rather than a protocol-dependent effect.
@@
Control measurements performed on distilled water (Figure 8a) show no oscillations, thresholds, or plateau structure, whereas DNA solutions at 1000 and 500 ng L-1 (Figs. 8b and 8c) exhibit the characteristic double-peak motif followed by discrete polarization plateaus. These features are reproducible in both magnetic-field sweep directions (arrows), indicating sweep-direction independence and the absence of hysteresis. Together, these comparisons establish that structured DNA hydration layers are required for the emergence of coherent polarization dynamics.
To assess the role of ionic contributions, we examined unbuffered DNA dissolved directly in distilled water (Figure 9). Both the 100 and 480 ng L-1 samples exhibit a pronounced threshold near T (voltage drop mV), followed by stabilization into discrete Landau-like polarization levels within the polarization-quantization formalism. The characteristic spacing between successive levels is approximately 1 mV, in close agreement with the previously reported value of mV. This agreement indicates that the observed quantization is intrinsic to the hydrated DNA matrix and does not depend on buffer ions. A Cu-coated control sample exhibits a transverse response of opposite sign, consistent with electronic conduction, further distinguishing the polarization-driven response observed in DNA solutions.
@@
Beyond the quantized steps and -periodic oscillations, coherent coupling between the longitudinal and transverse polarization channels is directly revealed by the interference reconstruction shown in Figure 10. By combining the simultaneously acquired signals and during a linear magnetic-field ramp, the interference map exhibits alternating constructive and destructive bands whose spacing and contrast remain stable across the entire 100–200 mT interval at 10 for 100 ng L-1 DNA. The region of maximal fringe intensity occurs near , consistent with the independently measured –phase lag and with Lissajous traces obtained under identical conditions (not shown). The persistence and sharpness of these interference fringes indicate that the two channels share a common, phase-coherent polarization mode. In this regime, the system behaves as a coupled dipolar oscillator with a well-defined phase relationship between its longitudinal and transverse components, reinforcing the interpretation that the transverse response originates from coherent polarization dynamics within the DNA–water hydrogen-bond network.
@@
The approximate magnetic-field–temperature (–) phase diagram shown in Figure 13 summarizes the regimes of coherent polarization identified across all measurements performed to date. The diagram incorporates magnetic-field thresholds (0.25–0.35 T), two reproducible temperature transitions at 20.6 and 12.0 , polarization staircases, bands of -periodic oscillations, and the emergence of low-temperature Fröhlich-type oscillations. Together, these observations delineate three principal domains: (i) a high-temperature incoherent regime without quantization; (ii) an intermediate regime in which discrete polarization plateaus coexist with -periodic oscillations; and (iii) a low-temperature, high-field regime characterized by a globally phase-locked polarization mode. Gaussian blurring was applied during construction to avoid artificially sharp boundaries, reflecting the intrinsic softness and heterogeneity of the hydrogen-bond network. The resulting diagram illustrates how magnetic alignment and thermal suppression of decoherence jointly organize the DNA–water matrix into quantized, and ultimately coherent, polarization states.
In addition to the transverse polarization response discussed above, the longitudinal transport channel exhibits a clear dynamical crossover. Under a fixed magnetic field and during gradual cooling, the initially ohmic longitudinal current collapses abruptly and subsequently fluctuates around zero, indicating a departure from steady dissipative conduction. Below this crossover, the longitudinal signal is dominated by low-amplitude, irregular fluctuations rather than a sustained DC component. This behavior is consistent with the emergence of a polarization-dominated regime, in which collective field reorganization replaces conventional charge transport as the dominant electrical response. A closely related longitudinal crossover, observed under comparable experimental conditions, has been reported in Ref. [18]. The coincidence of these observations supports the interpretation that the transverse response analyzed here arises in a regime where longitudinal conduction is suppressed, and polarization dynamics dominate.